Builtrite stock is currently selling for $40 a share. You purchase 200 shares and later the stock is selling for $52 a share.  What stock price would trigger a margin call (using 25%)?

Group of answer choices

$21.33

$42.67

$32.00

$64.00
Answer #1

Initial Deposit = 200 * 40 = $ 8000

We can solve this by looking at choices,

Margin call will occur when our deposit fall below the requirement margin. Hence, it will happen only when we are making loss.

Hence, any option greater than $40 won't result in margin call as it will be a profit to us.

Option $32.00

At $ 32.00

Amount with broker = 32 * 200 = 6400

Now, 6400 / 8000 = 80%

Hence, we have a loss of 20% only and hence no margin call

Option $ 21.33

At $ 21.33

Amount with broker = 21.33 * 200 = 4266

Now, 4266 / 8000 = 53.325%

Hence, we have a loss of 46.675% which is greater than 25% and hence there will be a margin call